(log12) X + (log12)(X-1) = 1
Answer by fractalier(6550)

Here we use the laws of logs to combine the logs at the left into one log:
log(12) [x(x-1)] = 1
Then we exponentiate both sides using 12 as our base:
12 ^ log(12) [x(x-1)] = 12 ^ 1
x(x-1) = 12
x^2 - x - 12 = 0
(x - 4)(x + 3) = 0
x = 4 and x = -3
Since we cannot take the log of a negative number,
only x = 4 is our solution.